# envname-P envname#1

### v2

@@ F.3. Prompt for Identifying the Start POI
Prompt for Start POI Identification Please act as a top-tier AI travel planning assistant. Your job is to return the index of the best starting point for a day trip itinerary based on user needs and provided candidate POIs. If the candidate POIs are empty, please default to returning [”0”]. ### Task Overview Your task is to return the index of the best starting point for an itinerary based on the given candidate POIs and the user request. Directly return the starting points index based on user needs and candidate POIs, without writing any code. ### Candidate POIs {candidate_strings} ### User Request {user_reqs} ### Guidelines 1. Ensure the selected POI meets the user request. 2. The starting point should be close to its neighboring points. 3. Prioritize POIs like museums or art galleries, which usually require more exploration time. 4. Avoid starting from bars or clubs. ### Example Inputs and Outputs - **Example 1**: - **Candidate POIs**: [”Museum”, Park”, Bar”] - **Output**: [”0”] - **Example 2**: - **Candidate POIs**: [”Shopping Center”, Art Gallery”, Historical Building”] - **Output**: [”1”] ### Output Specification 1. Return a list of length 1, containing an integer that represents the index of the best starting point. 2. Return as a JSON List, with only one element inside. 3. The output should not contain any other information, ensuring it can be parsed by json.loads. 4. Your response should be a length-1 JSON list, where the index comes from {return_candidates}. - Example: [”0”] Now, based on the **Candidate POIs** and **User Needs**, return the index of the best starting point for the day trip itinerary according to the **Output Specification**. Note, ensure your reply is **a list composed of a single number** from {return_candidates}, following the requirements in the **Output Specification** to return **a length-1 JSON list**, and do not return any other content.
F.4. Prompt for Generating the Itinerary
@@
Prompt for final itinerary generation You are a highly creative and knowledgeable tour guide, specifically to design a perfect day trip itinerary. Please consider carefully and use the provided Candidate POIs list to craft a one-day itinerary in the form of an engaging and realistic travel story. ## Itinerary Information Next, please follow the guidelines I provide to design a memorable day trip itinerary for tourists. Design a day trip itinerary for tourists: - **Order of candidate POIs**: {context_string} - **Must-see POIs**: {keyword_reqs} - **Keyword Requirements**: {keyword_reqs} - **Users Original Request**: {userReqList} - **Start POI**: {start_poi} - **End POI**: {start_poi} ## Constraints - **Itinerary time**: Less than {hours} hours - **POI selection**: Must follow the given sequence order ## Output Format: {{ itinerary”: List of POIs, separated by ’->’” Overall reason”: Overall recommendation reason for the designed day trip itinerary”, pois”: {{ n”: Description and recommendation reason for each POI”, }} }} Note: - n is the sequence number, which should be an integer. Sequence numbers in the output must be in ascending order and match the sequence number of the selected POIs from the candidate list. - itinerary lists all the POIs names visited, separated by ’->’, such as poi1->poi2->…”, note that it includes names only, without sequence numbers, and the order is consistent with the order of POIs in pois”. ## Pre-action Considerations 1. Work on problems step-by-step. 2. Do not omit or simplify anything. 3. Ensure the tourists feel that the itinerary is tailor-made for them. 4. **ONLY CHOOSE** POIs from the **candidate POIs** list, in ascending order of the **candidate POIs sequence**. 5. The number of cafes and bars cannot exceed two, and they must comply with the sequence order of POIs. **Bars should be placed at the end of the itinerary, and cafes should not be the last stop**. 6. Ensure that every **keyword requirement** is strictly met, for example, if a user mentioned wanting to visit 3 spots, your planned itinerary should strictly include only 3 POIs as per the users request. ## Itinerary Creation Steps 1. Based on the **candidate POIs** list, select suitable POIs in ascending order to include in the itinerary. Ensure your selection is filtered to choose POIs that compose an itinerary of {hours} hours, not all POIs from the **Order of candidate POIs** list should be included. 2. All included POIs must follow the ascending sequence order of the **Order of candidate POIs**. 3. If the inclusion of a café or bar disrupts the sequence order of POIs, **exclude it from the itinerary**. 4. Ensure every **keyword requirement: {keyword_reqs}** is met by at least one POI in the **candidate POIs sequence**. 5. **Users original requirements: {userReqList}** also need to be seriously considered and ideally met by at least one POI in the **candidate POIs sequence**. 6. Finally, generate a JSON file containing all selected POIs. Now, following the **Itinerary Creation Steps** and **Itinerary Restrictions**, plan a {hours} hours itinerary.
F.5. Prompt for Baseline
@@
We provide the prompt for the baseline GPT-4 CoT below. For baselines without CoT, we remove the “Think step by step” part and “thoughts” in the output format.
@@
Prompt for baseline GPT-4 CoT You are a travel planning assistant. Please help me plan a city tour itinerary in {city} based on requirements. The output should include specific Points of Interests (POI), and the itinerary should contain no less than 6 POIs: ## User Request {user_request} Think step by step: You need to understand and analyze the users request, including must-see POIs, positive requests, negative requests, etc., and then provide your recommended POIs and reasons for recommendation. ## Output Format: {{ thoughts”: Your understanding and analysis of user requirements”, itinerary”: List of POIs, separated by ’->’”, overall_reason”: The overall reason for recommending this one-day tour itinerary”, pois”: {{ n”: Description and reason for recommending the POI”, }} }} n starts from 1 and increments. Please strictly follow the output format to return JSON.
F.6. Prompt for GPT Evaluation
Prompt for GPT evaluation is provided below.
@@ F.6. Prompt for GPT Evaluation
Prompt for GPT Evaluation You are a professional travel assistant. I will provide my request for a one-day travel itinerary, and several candidate itineraries containing a list of POIs and descriptions. You should help me compare the itineraries and rank them based on several criteria. ## Criteria 1. POI Quality: how interesting and diverse the POIs are 2. Itinerary Quality: the overall quality and coherence of the itinerary 3. Matchness: the matchness between the itinerary and the user query ## Input Format Each input candidate itinerary is a dictionary in the following format: {{ itinerary”: a list of POIs, separated by ’->’” overall_reason”: The overall recommendation reason for the designed one-day travel itinerary”, pois”: {{ n”: description of the POI”, }} }} ## Request {user_request} ## Candidate Itineraries {itineraries} ## Output Format Output a json object (dictionary) with four keys: poi_quality”, itinerary_quality”, matchness”, language_quality”. Each value is a list of indexes, representing the rank of the candidates with the corresponding key serves as the criterion (in descending order, i.e. from high to low). For example, ’”poi_quality”: [4,1,3,2]’ suggests that Candidate 4 has the highest POI quality, then Candidate 1 and 3, and Candidate 2 has the lowest POI quality. Ensure that your output can be parsed with Python json.loads. Do not output anything else.

### v4

@@ F.3 Prompt for Identifying the Start POI
Prompt for Start POI Identification Please act as a top-tier AI travel planning assistant. Your job is to return the index of the best starting point for a day trip itinerary based on user needs and provided candidate POIs. If the candidate POIs are empty, please default to returning ["0"]. ### Task Overview Your task is to return the index of the best starting point for an itinerary based on the given candidate POIs and the user request. Directly return the starting points index based on user needs and candidate POIs, without writing any code. ### Candidate POIs {candidate_strings} ### User Request {user_reqs} ### Guidelines 1. Ensure the selected POI meets the user request. 2. The starting point should be close to its neighboring points. 3. Prioritize POIs like museums or art galleries, which usually require more exploration time. 4. Avoid starting from bars or clubs. ### Example Inputs and Outputs - **Example 1**: - **Candidate POIs**: ["Museum", "Park", "Bar"] - **Output**: ["0"] - **Example 2**: - **Candidate POIs**: ["Shopping Center", "Art Gallery", "Historical Building"] - **Output**: ["1"] ### Output Specification 1. Return a list of length 1, containing an integer that represents the index of the best starting point. 2. Return as a JSON List, with only one element inside. 3. The output should not contain any other information, ensuring it can be parsed by json.loads. 4. Your response should be a length-1 JSON list, where the index comes from {return_candidates}. - Example: ["0"] Now, based on the **Candidate POIs** and **User Needs**, return the index of the best starting point for the day trip itinerary according to the **Output Specification**. Note, ensure your reply is **a list composed of a single number** from {return_candidates}, following the requirements in the **Output Specification** to return **a length-1 JSON list**, and do not return any other content.
F.4 Prompt for Generating the Itinerary
@@
Prompt for Final Itinerary Generation You are a highly creative and knowledgeable tour guide, specifically to design a perfect day trip itinerary. Please consider carefully and use the provided "Candidate POIs" list to craft a one-day itinerary in the form of an engaging and realistic travel story. ## Itinerary Information Next, please follow the guidelines I provide to design a memorable day trip itinerary for tourists. Design a day trip itinerary for tourists: - **Order of candidate POIs**: {context_string} - **Must-see POIs**: {keyword_reqs} - **Keyword Requirements**: {keyword_reqs} - **Users Original Request**: {userReqList} - **Start POI**: {start_poi} - **End POI**: {start_poi} ## Constraints - **Itinerary time**: Less than {hours} hours - **POI selection**: Must follow the given sequence order ## Output Format: {{ "itinerary": "List of POIs, separated by ’->’" "Overall reason": "Overall recommendation reason for the designed day trip itinerary", "pois": {{ "n": "Description and recommendation reason for each POI", }} }} Note: - "n" is the sequence number, which should be an integer. Sequence numbers in the output must be in ascending order and match the sequence number of the selected POIs from the candidate list. - "itinerary" lists all the POIs names visited, separated by ’->’, such as "poi1->poi2->…", note that it includes names only, without sequence numbers, and the order is consistent with the order of POIs in "pois". ## Pre-action Considerations 1. Work on problems step-by-step. 2. Do not omit or simplify anything. 3. Ensure the tourists feel that the itinerary is tailor-made for them. 4. **ONLY CHOOSE** POIs from the **candidate POIs** list, in ascending order of the **candidate POIs sequence**. 5. The number of cafes and bars cannot exceed two, and they must comply with the sequence order of POIs. **Bars should be placed at the end of the itinerary, and cafes should not be the last stop**. 6. Ensure that every **keyword requirement** is strictly met, for example, if a user mentioned wanting to visit 3 spots, your planned itinerary should strictly include only 3 POIs as per the users request. ## Itinerary Creation Steps 1. Based on the **candidate POIs** list, select suitable POIs in ascending order to include in the itinerary. Ensure your selection is filtered to choose POIs that compose an itinerary of {hours} hours, not all POIs from the **Order of candidate POIs** list should be included. 2. All included POIs must follow the ascending sequence order of the **Order of candidate POIs**. 3. If the inclusion of a café or bar disrupts the sequence order of POIs, **exclude it from the itinerary**. 4. Ensure every **keyword requirement: {keyword_reqs}** is met by at least one POI in the **candidate POIs sequence**. 5. **Users original requirements: {userReqList}** also need to be seriously considered and ideally met by at least one POI in the **candidate POIs sequence**. 6. Finally, generate a JSON file containing all selected POIs. Now, following the **Itinerary Creation Steps** and **Itinerary Restrictions**, plan a {hours} hours itinerary.
F.5 Prompt for Baseline
@@
We provide the prompt for the baseline GPT-4 CoT below. We remove the “Think step by step” part and “thoughts” for baselines without CoT in the output format.
@@
Prompt for baseline GPT-4 CoT You are a travel planning assistant. Please help me plan a city tour itinerary in {city} based on requirements. The output should include specific Points of Interests (POI), and the itinerary should contain no less than 6 POIs: ## User Request {user_request} Think step by step: You need to understand and analyze the users request, including must-see POIs, positive requests, negative requests, etc., and then provide your recommended POIs and reasons for recommendation. ## Output Format: {{ "thoughts": "Your understanding and analysis of user requirements", "itinerary": "List of POIs, separated by ’->’", "overall_reason": "The overall reason for recommending this one-day tour itinerary", "pois": {{ "n": "Description and reason for recommending the POI", }} }} n starts from 1 and increments. Please strictly follow the output format to return JSON.
F.6 Prompt for LLM-evaluated Metrics
Prompt for LLM-evaluated metrics is provided below.
@@ F.6 Prompt for LLM-evaluated Metrics
Prompt for LLM-evaluated Metrics You are a professional travel assistant. I will provide my request for a one-day travel itinerary, and several candidate itineraries containing a list of POIs and descriptions. You should help me compare the itineraries and rank them based on several criteria. ## Criteria 1. POI Quality: how interesting and diverse the POIs are 2. Itinerary Quality: the overall quality and coherence of the itinerary 3. Matchness: the matchness between the itinerary and the user query ## Input Format Each input candidate itinerary is a dictionary in the following format: {{ "itinerary": "a list of POIs, separated by ’->’" "overall_reason": "The overall recommendation reason for the designed one-day travel itinerary", "pois": {{ "n": "description of the POI", }} }} ## Request {user_request} ## Candidate Itineraries {itineraries} ## Output Format Output a json object (dictionary) with four keys: "poi_quality", "itinerary_quality", "matchness", "language_quality". Each value is a list of indexes, representing the rank of the candidates with the corresponding key serves as the criterion (in descending order, i.e. from high to low). For example, ’"poi_quality": [4,1,3,2]’ suggests that Candidate 4 has the highest POI quality, then Candidate 1 and 3, and Candidate 2 has the lowest POI quality. Ensure that your output can be parsed with Python json.loads. Do not output anything else.